# ATOMIUM: Dust and tracers of binarity in the continua

### v1

@@ 4.5 Trends with period
with the RSGs found above this trend line. If we assume that the Mira variables are all fundamental mode pulsators, this is in general agreement with the theoretical models of Ahmad et al. (2023) who found an increasing period with radius based on their CO5BOLD simulations.
@@
However, for the majority of the sample, we do not see any clear trends beyond what was already apparent from Fig. 4: the SR AGB stars have lower fluxes than the Miras. The Miras with periods days all have comparable monochromatic luminosities with no correlation to the period. This is quite different to period-luminosity diagrams constructed from bolometric luminosities or K-band absolute magnitudes (e.g. Whitelock et al., 2008; Trabucchi et al., 2017), which show much clearer trends for both Miras and SR variables.
@@ 5.1 Dust features
With these observations, we present for the first time millimetre stellar and dust continuum emission at high angular resolution for a sizeable sample of AGB and RSG stars. Where our ALMA continuum maps have been discussed phenomenologically in Sect. 3, we now focus on explanations for the observed dust distributions, including some comparisons between the ALMA continuum and SPHERE observations. Part of the ATOMIUM campaign involved contemporaneous (within 10 days) observations using the ALMA extended configuration and the Very Large Telescope’s Spectropolarimetric High-contrast Exoplanet Research instrument with the Zurich Imaging Polarimeter (VLT/SPHERE-ZIMPOL, Beuzit et al., 2019; Schmid et al., 2018, hereafter referred to as SPHERE). These have been published in detail by Montargès et al. (2023) and here we pay particular attention to the degree of linear polarisation (DoLP) plots in their Figure 3.
@@ 5.1.1 Gru
This tail of dust lies in the wake of the C companion of the Gru system. By collecting archival SPHERE data, Montargès et al. (2025) have shown that the tail changes position as the companion orbits the AGB star. Subsequent ALMA observations have confirmed that the continuum feature associated with the C component also changes position in an orbit around the AGB star (Esseldeurs et al, subm.). Using hydrodynamic simulations, El Mellah et al. (2020) and Malfait et al. (2024a) showed that not only do close companions to AGB stars shape the AGB wind into spiral patterns (especially when viewed face on, with the line of sight perpendicular to the orbital plane) but that accretion discs can also form around such companions. Montargès et al. (2025) presented a detailed analysis of the nature of the C companion, and concluded that the feature attributed to the C companion in the extended ALMA continuum likely originates from an accretion disc around the star (which may be a white dwarf of a K dwarf main sequence star).
Montargès et al. (2025) suggest that the dust tail is seen because dust is concentrated in the wake of the companion or that dust growth is accelerated in that region. Based on the observations presented here of Gru and other stars (see discussion on W Aql below) we suggest that dust preferentially forms in the wakes of shocks caused by companions moving supersonically through the AGB wind. Since the sound speed is relatively low, even in the inner wind ( km s-1), the orbital velocity will almost always be supersonic for a companion within the AGB circumstellar envelope. This is the origin of the shock waves seen in hydrodynamic simulations of companion-wind interactions (e.g. Maes et al., 2021; Malfait et al., 2021, 2024a). These shocked structures are denser than the surrounding gas, meaning that chemical reactions, including those involved in dust formation, can progress faster. This would also explain the dust distribution around IK Tau (see discussion in Coenegrachts et al. 2023) and is in line with other theoretical modelling results which show that dust forms in the wake of shocks (Freytag & Höfner, 2023). We ran a hydrodynamic model including dust formation to test this point and present some first results in Appendix F. A detailed theoretical study of binary-enhanced dust formation over a wider parameter space is forthcoming (Samaratunge et al, in prep).
@@ 5.1.2 W Aql
W Aql has a known F9 main sequence companion presently at a separation of around 200 au (see detailed discussion in Danilovich et al., 2024). In Fig. LABEL:fig:waql, continuum emission is detected close to the position of the companion with a confidence of . Since we know the nature of the F9 companion, we know that the star itself cannot account for that level of 1.24 mm emission. The hydrodynamic models in Danilovich et al. (2024) do not predict an accretion disc around such a distant companion888Though we note this could be because those models do not include cooling, see discussion in Malfait et al. (2024a). (see also El Mellah et al., 2020), but they do predict a shocked wake close to the companion. The dust observed near the companion could hence be formed in its wake. If this is the case, then we are observing some dust formation au from the AGB star, much further out than the dust condensation radius, estimated to be around 13 au by Danilovich et al. (2014).
@@ 5.1.3 GY Aql
The bar seen in the GY Aql continuum images is found at a larger separation from the position of the star than the DoLP SPHERE image presented in Montargès et al. (2023), where the majority of the DoLP flux is within . The SPHERE image shows arcs of polarised flux around the star, especially to the south and northwest. Figure 11 of Montargès et al. (2023) shows that the DoLP image may trace out some of the spiral pattern seen in the CO moment 1 map, which represents the velocity field. Since the CO channel maps reveal spiral-like patterns in the wind of GY Aql (Decin et al., 2020), it is not unreasonable to expect (some of) the dust to also follow the density structures in the wind. Unfortunately, at the separation of the bar structure from the AGB star, the velocity field is messy and it is difficult to draw conclusions through a simple comparison. A more detailed analysis of GY Aql is forthcoming in a separate paper (Marinho et al, in prep).
@@ 5.1.4 SV Aqr
The continuum emission observed with ALMA towards SV Aqr (Fig. LABEL:fig:svaqr) appears to form an arc. The DoLP image presented in Montargès et al. (2023) has rather low S/N and only a small region of polarised flux above the detection threshold, located to the southeast of the star and within mas. This is too close to be resolved by the ALMA beam in the extended data, but is close to the asymmetry seen in the contour of the extended data.
Overall, the dust in the vicinity of SV Aqr follows an irregular pattern. It could be distributed in this way as a result of episodic dust formation, perhaps driven by asymmetric bubbles on the surface, similar to what is seen in the models of Freytag & Höfner (2023). If the dust emission truly forms an arc or loop near the star, then this shape could possibly have been formed by a companion shaping the wind and/or enhancing dust in this region. We found SV Aqr to have an additional period of 231.8 days, with alternating cycles having minima of different depths (see Appendix A for details). This could indicate a binary system with a period of 463.6 days. However, the feasibility of this strongly depends on the system mass. If the system mass is only (i.e. for a planetary or brown dwarf companion), then the separation for a circular orbit would be only 1.2 au, smaller than the radius we measured with ALMA, even considering our significant uncertainties. A system mass of or higher would place a companion (just) outside of our measured radius.
@@ 5.1.6 KW Sgr
The continuum images of KW Sgr do not have notable features, aside from a protrusion to the east-southeast of the UD subtracted image (Fig. LABEL:fig:kwsgr). Although this protrusion has a low S/N, it agrees well with a similar protrusion seen in the SPHERE DoLP presented by Montargès et al. (2023), although the DoLP feature also has a low S/N. To emphasise their similarity, we plot the UD-subtracted continuum contours over the SPHERE DoLP image in Fig. 7. The fact that two very different imaging techniques show a similar protrusion, lends some credence to it. Given the high asymmetry of this feature, we suggest that it may be a result of episodic mass loss, which is thought to be characteristic of RSGs (e.g. see Levesque 2017 and references therein, the phenomenon termed the Great Dimming of Betelgeuse, Montargès et al. 2021, and the episodic ejecta of VY CMa, Humphreys et al. 2024, 2025).
@@ 5.1.7 VX Sgr and AH Sco
In the SPHERE observations presented in Montargès et al. (2023), the RSGs VX Sgr and AH Sco exhibit moderately extended regions of polarised flux. In both cases, especially AH Sco, the highlighted regions are partly limited by the S/N of the observations. That is, we see polarised flux mainly to the north and south of AH Sco with SPHERE but this is partly because the S/N to the east and west is worse, limiting our ability to compare with the significant ALMA flux to the east and west of AH Sco (Sect. 3.4.5). For VX Sgr, there appears to be more polarised flux from SPHERE than extended emission from ALMA, but this is likely a result of ALMA resolving out some flux (see Sect. 3.3.3). The VLT/AMBER imaging of VX Sgr in the near infrared by Chiavassa et al. (2010) also shows significant complexity in the atmosphere and within a radius of mas. Chiavassa et al. (2010) tentatively show extended structures close to the star ( mas) to the north and south. Our extended array observations (Fig. LABEL:fig:vxsgr) show structures to the north and south out to mas. If these are the same ejecta and we consider the difference in observation times ( 11 years), then the lower limit on the velocity of these features (ignoring projection effects) would be 60 km s-1, comparable to some of the high velocity jets seen for VY CMa, another extreme RSG (Quintana-Lacaci et al., 2023). If these are not the same structures, then their similarity across physical scales at different times are suggestive of localised episodic mass loss around VX Sgr. What we can conclude from the infrared and ALMA observations is that both VX Sgr and AH Sco have significant circumstellar dust, as has been shown by earlier observations of, for example, silicate features in infrared spectra (Speck et al., 2000).
@@ 5.2 Angular sizes at millimetre wavelengths
Prior to the present work, a few studies have examined the angular sizes of AGB stars at sub-mm, mm and longer radio wavelengths. To ensure a valid comparison, we focus on the observations closest to the central frequency of our Band 6 continua (241.75 GHz or 1.24 mm). Vlemmings et al. (2019) studied four nearby oxygen-rich AGB stars (W Hya, Cet (aka Mira A), R Dor, and R Leo) at high angular resolution across multiple wavelength ranges. Higher resolution observations of W Hya (Ohnaka et al., 2024) and R Dor (Vlemmings et al., 2024) are also available. Uniform discs were fit by the authors of these earlier observations, allowing us to directly compare the stellar properties to our sample.
Of these four stars, R Dor is a SRb type variable, while W Hya, Cet and R Leo are all Mira variables. Situating them in our plot of monochromatic luminosity against UD radius in Fig. 4, we find that R Dor sits in the SR quadrant of the plot, with a lower flux and smaller radius than the Mira variables. The three Mira stars lie in the Mira quadrant, upholding the divisions we found. When considering the pulsation periods of these stars and comparing with our plots of UD properties against period in Fig. 5, we find that the Mira variables cluster close to the Miras in our sample (with periods of 402 days, 332 days and 310 days for W Hya, Cet, and R Leo, respectively). R Dor has two periods of comparable length to T Mic (362 and 175 days for R Dor, Vlemmings et al., 2024, 352 and 178 days for T Mic, Table 1), so sits near T Mic and S Pav on the plot of UD radius against period if we consider the longer period as the primary.
@@
The recent work of Vlemmings et al. (2024) measured the stellar disc of the closest AGB star, R Dor, with ALMA and compared this diameter with the NIR diameter measured by Ohnaka et al. (2019). For the ALMA radius at 225 GHz (1.33 mm, close to our 1.24 mm) the measurements of Vlemmings et al. (2024) and Ohnaka et al. (2019) agree well with our Eq. 2. They report for R Dor , which is the same gradient as in our Eq. 2, but with an intercept of 0,0. As mathematically expected, a line with the same slope passing through 0,0 does not agree with our data as well as Eq. 2. If we instead fit to our data but force the line to pass through 0,0, we find a slope of 1.38, which is a reasonable fit to the measured data. Of course our Eq. 2 must only be valid over a limited regime, since a non-zero diameter at one wavelength cannot physically correspond to a diameter of zero in another. However, in the right panel of Fig. 3 we plot observed sizes not physical sizes and the real lower limits on both axes are the resolution limits. Fitting to physical sizes is not possible for our dataset because most of the measured stars have very similar radii (see Table 3 and Fig. 4).
@@
5.3 R Aql and R Hya: recent thermal pulses?
If R Aql and R Hya both underwent a relatively recent (past few hundred years) thermal pulse, as suggested by Wood & Zarro (1981), Greaves & Howarth (2000) and Zijlstra et al. (2002), then we would perhaps expect to see some similarities in their circumstellar envelopes. This is not true of their continuum images (R Aql: Fig. 2; R Hya: Fig. LABEL:fig:rhya), where R Aql has a relatively featureless continuum, while R Hya has a lot of extended circumstellar dust. While it is possible that R Aql suffers from more resolved out flux, it is a little over twice as far away, suggesting a smaller angular size (all else being equal) and hence being less prone to resolved out flux. The ALMA continuum observations are in general agreement with the SPHERE observations of Montargès et al. (2023), which reveal a larger region of polarised flux for R Hya than for R Aql.
Despite their difference in dust, R Aql and R Hya do have some similarities. The physical radii derived from our UD fits and the monochromatic luminosities for both stars are close to being the same, within our uncertainties. They are also the two Mira stars with the smallest radii when comparing physical units, both from our UD fits (Table 2) and from literature measurements taken in the NIR (Table 3). From their study of dust in the infrared, Zhao-Geisler et al. (2012) estimated that the inner edges of the IR dust shells around these stars are au and au from the corresponding stars, for R Aql and R Hya respectively. Strikingly, R Aql shows an increase in flux in the far infrared (in the ISO LWS spectrum, see Figure 2 of Zhao-Geisler et al. 2012), suggesting a shell of cool dust, as might be expected for a detached dust shell. The same is not seen for R Hya (see also the earlier results of Decin et al., 2008). If R Aql truly has a burgeoning detached shell while R Hya — which has a significant amount of dust close to the star — does not, that would imply that R Hya had a more recent thermal pulse. This is counter to the analysis based on luminosity done by Wood & Zarro (1981), which predicts that less time has passed since the thermal pulse of R Aql than R Hya. However, Maercker et al. (2016, 2024) presented ALMA observations that demonstrate that for some of the surveyed (carbon-rich) AGB stars the shells thought to have been formed through thermal pulses (1000–8000 years ago) were followed by a moderate decrease in mass loss, rather than an abrupt decrease seen in other cases. It is possible that R Hya has undergone a more moderate decrease in mass-loss post-thermal pulse, while R Aql has undergone a more drastic change.
On larger scales than our ALMA observations, a dust shell around R Hya was proposed by Hashimoto & Izumiura (1998) based on IRAS spectroscopy and photometric imaging. Higher resolution Herschel/PACS imaging reveals a fermata-style bow shock around R Hya (Cox et al., 2012), which Maercker et al. (2022) modelled as a partial ellipsoid shell approximately 100″from the AGB star. This more extended dust cannot be captured in our ALMA observations, partly because much of it falls outside of our field of view, and partly because large scale structures tend to be resolved out by ALMA. Similar observations of R Aql are not available; R Aql was not imaged with Herschel/PACS.
The CO channel maps of of R Aql and R Hya, presented in Decin et al. (2020), show unique patterns for both stars, dissimilar to each other and to the other stars in the ATOMIUM sample. R Aql has a rose-like pattern, which Malfait et al. (2024b) was unable to reproduce through a hydrodynamical model of a hierarchical triple star system999Malfait et al. (2024a) found that a qualitatively similar pattern seen in the earlier hydrodynamic models of El Mellah et al. (2020) and Malfait et al. (2021) was most likely caused by periodic instabilities arising from an absence of cooling functions aside from adiabatic expansion.. A possibility is that the rose-like structure was caused by a thermal pulse and that the irregular arc-like pattern is an imprint of the convective flows in the stellar atmosphere (compare with the gas densities in the models of Freytag & Höfner 2023 and the irregularities in the otherwise thin detached shells seen by Maercker et al. 2024). From the angular extent of the shapes in the CO emission, the distance to the star, and the expansion velocities (Decin et al., 2020; Wallström et al., 2024), this would imply a thermal pulse –1000 years ago.
R Hya has at least three pairs of bipolar structures, as described in detail by Homan et al. (2021). Decin et al. (2008) estimated a recent decrease in mass loss by a factor of 20 for R Hya, which they attribute to a thermal pulse period ending 220 years ago, in agreement with the timeline proposed by Wood & Zarro (1981). However, it is unclear how the complicated circumstellar structure revealed by the ATOMIUM molecular lines (Homan et al., 2021) and the dust in the continuum observations (Fig. LABEL:fig:rhya) can be reconciled with a recent thermal pulse. A possible explanation could be a companion star in the inner wind, disrupting the more spherical shell that is expected to form during a thermal pulse.

### v2

@@ 4.5 Trends with period
with the RSGs found slightly above this trend line. If we assume that the Mira variables are all fundamental mode pulsators, this is in general agreement with the theoretical models of Ahmad et al. (2023) who found an increasing period with radius based on their CO5BOLD simulations.
@@
However, for the majority of the sample, we do not see any clear trends beyond what was already apparent from Fig. 6: the SR AGB stars have lower fluxes than the Miras. The Miras with periods days all have comparable monochromatic luminosities with no correlation to the period. There is a weak positive correlation if we consider both Miras with periods days and SRs, but this arises from the fact that the SRs in our sample generally have shorter periods and all have lower monochromatic luminosities than the Miras. Fig. 7 is quite different to period-luminosity diagrams constructed from bolometric luminosities or K-band absolute magnitudes (e.g. Whitelock et al., 2008; Trabucchi et al., 2017), which show much clearer trends for both Miras and SR variables.
@@ 5.1 Dust features
With these observations, we present for the first time millimetre stellar and dust continuum emission at high angular resolution for a sizeable sample of AGB and RSG stars. Where our ALMA continuum maps have been discussed phenomenologically in Sect. 3, we now focus on explanations for the observed dust distributions, including some comparisons between the ALMA continuum and observations from the Very Large Telescope’s Spectropolarimetric High-contrast Exoplanet Research instrument with the Zurich Imaging Polarimeter (VLT/SPHERE-ZIMPOL, Beuzit et al., 2019; Schmid et al., 2018, hereafter referred to as SPHERE). Part of the ATOMIUM campaign involved contemporaneous (within 10 days) observations using the ALMA extended configuration and SPHERE. These have been published in detail by Montargès et al. (2023) and here we pay particular attention to the degree of linear polarisation (DoLP) plots in their Figure 3. Note that the SPHERE images mainly show scattered light from dust grains located in the plane of the sky (see discussion in Montargès et al., 2023), while our ALMA continuum images recover thermal emission from dust grains, in addition to the contribution to the continuum flux from the weak stellar photosphere/chromosphere and the radio photosphere as first described in Reid & Menten (1997) — see Sect. 5.2.
@@ 5.1.1 Gru
This tail of dust lies in the wake of the C companion of the Gru system. By collecting archival SPHERE data, Montargès et al. (2025) have shown that the tail changes position as the companion orbits the AGB star with a period of 11 years. Subsequent ALMA observations have confirmed that the continuum feature associated with the C component also changes position in an orbit around the AGB star (Esseldeurs et al, subm.). Using hydrodynamic simulations, El Mellah et al. (2020) and Malfait et al. (2024a) showed that not only do close companions to AGB stars shape the AGB wind into spiral patterns (especially when viewed face on, with the line of sight perpendicular to the orbital plane) but that accretion discs can also form around such companions. Montargès et al. (2025) presented a detailed analysis of the nature of the C companion, and concluded that the feature attributed to the C companion in the extended ALMA continuum likely originates from an accretion disc around the star (which may be a white dwarf of a K dwarf main sequence star).
Montargès et al. (2025) suggest that the dust tail is seen because dust is concentrated in the wake of the companion or that dust growth is accelerated in that region. Based on the observations presented here of Gru and other stars (see discussion on W Aql below) we suggest that dust preferentially forms in the wakes of shocks caused by companions moving supersonically through the AGB wind. Since the sound speed is relatively low, even in the inner wind ( km s-1), the orbital velocity will almost always be supersonic for a companion within the AGB circumstellar envelope. This is the origin of the shock waves seen in hydrodynamic simulations of companion-wind interactions (e.g. Maes et al., 2021; Malfait et al., 2021, 2024a). These shocked structures are denser than the surrounding gas, meaning that chemical reactions, including those involved in dust formation, can progress faster. This would also explain the dust distribution around IK Tau (see discussion in Coenegrachts et al. 2023) and is in line with other theoretical modelling results which show that dust forms in the wake of shocks (Freytag & Höfner, 2023). We ran a hydrodynamic model including dust formation to test this point and present some first results in Appendix H. A detailed theoretical study of binary-enhanced dust formation over a wider parameter space is forthcoming (Samaratunge et al, in prep).
@@ 5.1.2 W Aql
W Aql has a known F9 main sequence companion presently at a separation of around 200 au (see detailed discussion in Danilovich et al., 2024). In Fig. LABEL:fig:waql, continuum emission is detected close to the position of the companion with a confidence of . Since we know the nature of the F9 companion, we know that the star itself cannot account for that level of 1.24 mm emission. The hydrodynamic models in Danilovich et al. (2024) do not predict an accretion disc around such a distant companion101010Though we note this could be because those models do not include cooling, see discussion in Malfait et al. (2024a). (see also El Mellah et al., 2020), but they do predict a shocked wake close to the companion. The dust observed near the companion could hence be formed in its wake. If this is the case, then we are observing some dust formation au from the AGB star. This is much further out than the dust condensation radius, estimated to be around 13 au by Danilovich et al. (2014), which is the approximate region within which all dust was previously thought to form.
@@ 5.1.3 GY Aql
The bar seen in the GY Aql continuum images (Fig. 2) is found at a larger separation from the position of the star (extending out to ″) than the DoLP SPHERE image presented in Montargès et al. (2023), where the majority of the DoLP flux is within . The SPHERE image shows arcs of polarised flux around the star, especially to the south and northwest. Figure 11 of Montargès et al. (2023) shows that the DoLP image may trace out some of the spiral pattern seen in the CO moment 1 map, which represents the velocity field. Since the CO channel maps reveal spiral-like patterns in the wind of GY Aql (Decin et al., 2020), it is not unreasonable to expect (some of) the dust to also follow the density structures in the wind. Unfortunately, at the separation of the bar structure from the AGB star, the velocity field is messy and it is difficult to draw conclusions through a simple comparison. A more detailed analysis of GY Aql is forthcoming in a separate paper (Marinho et al, in prep).
@@ 5.1.4 SV Aqr
The continuum emission observed with ALMA towards SV Aqr (Fig. LABEL:fig:svaqr) appears to form an arc. The DoLP image presented in Montargès et al. (2023) has rather low S/N and only a small region of polarised flux above the detection threshold, located to the southeast of the star and within mas. This is too close to be resolved by the ALMA beam in the extended data, but is close to the asymmetry seen in the contour of the extended data. We reproduce the DoLP image over-plotted with the extended continuum contours in Fig. 9, but we caution against over-interpretation of this image, and direct the reader to the discussion of reliability in Montargès et al. (2023) and the contours plotted in their Fig. 3.
Overall, the dust in the vicinity of SV Aqr follows an irregular pattern. It could be distributed in this way as a result of episodic dust formation, perhaps driven by asymmetric bubbles on the surface, similar to what is seen in the models of Freytag & Höfner (2023). Dust emission forming an arc or loop near the star could have been formed by a companion shaping the wind and/or enhancing dust in this region. We found SV Aqr to have an additional period of 231.8 days, with alternating cycles having minima of different depths (see Appendix A.1 for details). This could indicate a binary system with a period of 463.6 days. However, the feasibility of this strongly depends on the system mass. If the system mass is only (i.e. for a planetary or brown dwarf companion), then the separation for a circular orbit would be only 1.2 au, smaller than the radius we measured with ALMA, even considering our significant uncertainties. A system mass of or higher would place a companion (just) outside of our measured radius. Hence, the available evidence suggests a rather tight binary system. Since it is not presently interacting strongly, this further indicates a system mass .
@@ 5.1.6 KW Sgr
KW Sgr, a RSG that we did not observe with the compact configuration of ALMA, has a highly elliptical beam for the mid data and some small asymmetries in the continuum at levels of but generally smaller than the area of the beam, and hence not significant. The UD subtracted image (Fig. LABEL:fig:kwsgr) shows some extended dust to the east south-east of the continuum peak, with a confidence of , and a little more emission at the position of the continuum peak. Although this protrusion has a low S/N, it agrees well with a similar protrusion seen in the SPHERE DoLP presented by Montargès et al. (2023), although the DoLP feature also has a low S/N. To emphasise their similarity, we plot the UD-subtracted continuum contours over the SPHERE DoLP image in Fig. 10. The fact that two very different imaging techniques show a similar protrusion, lends some credence to it. Given the high asymmetry of this feature, we suggest that it may be a result of episodic mass loss, which is thought to be characteristic of RSGs (e.g. see Levesque 2017 and references therein, the phenomenon termed the Great Dimming of Betelgeuse, Montargès et al. 2021, and the episodic ejecta of VY CMa, Humphreys et al. 2024, 2025).
@@ 5.1.7 VX Sgr and AH Sco
In the SPHERE observations presented in Montargès et al. (2023), the RSGs VX Sgr and AH Sco exhibit moderately extended regions of polarised flux. In both cases, especially AH Sco, the highlighted regions are partly limited by the S/N of the observations. That is, we see polarised flux mainly to the north and south of AH Sco with SPHERE but this is partly because the S/N to the east and west is worse, limiting our ability to compare with the significant ALMA flux to the east and west of AH Sco (Fig. LABEL:fig:ahsco, Sect. 3.5.6). For VX Sgr, there appears to be more polarised flux from SPHERE than extended emission from ALMA, but this is likely a result of ALMA resolving out some flux (see Sect. 3.4.3). The VLT/AMBER imaging of VX Sgr in the near infrared by Chiavassa et al. (2010) also shows significant complexity in the atmosphere and within a radius of mas. Chiavassa et al. (2010) tentatively show extended structures close to the star ( mas) to the north and south. Our extended array observations (Fig. LABEL:fig:vxsgr) show structures to the north and south out to mas ( au). If these are the same ejecta and we consider the difference in observation times ( 11 years), then the lower limit on the velocity of these features (ignoring projection effects) would be 60 km s-1, comparable to some of the high velocity jets seen for VY CMa, another extreme RSG (Quintana-Lacaci et al., 2023). If these are not the same structures, then their similarity across physical scales at different times are suggestive of localised episodic mass loss around VX Sgr. What we can conclude from the infrared and ALMA observations is that both VX Sgr and AH Sco have significant circumstellar dust, as has been shown by earlier observations of, for example, silicate features in infrared spectra (Speck et al., 2000).
@@ 5.2 Angular sizes at millimetre wavelengths
| Star | Distance | Distance | Variability | Period | Period | UD diam. | Flux | at freq. | UD |
| [pc] | ref. | type | [days] | ref. | [mas] | [mJy] | [GHz] | ref. | |
| R Dor | 1 | SRb | 332, 175 | 5 | 61.8 | 221.4 | 225 | 1 | |
| W Hya | 2 | Mira | 402 | VSX | 50 a | 210 a | 240 | 3 | |
| Mira A | 3 | Mira | 331 | VSX | 43.3 | 99.61 | 229.13 | 3 | |
| R Leo | 2 | Mira | 312 | VSX | 41.88 | 107.83 | 232.53 | 3 | |
| R Lep | 2 | Mira | 437 | VSX | 14.2 | 37.3 | 405.0 | 6 | |
| CW Leo | 4 | Mira | 630 | VSX | 53 b | 372 | 258 c | 7 |
The apparent angular sizes of AGB stars at radio wavelengths are thought to be dominated by electron-neutral free-free emission out to (Reid & Menten, 1997; Matthews et al., 2015; Vlemmings et al., 2019). The apparent size of this radio photosphere decreases with increasing frequencies, because of changes in opacity, resulting in apparent sizes for GHz (Vlemmings et al., 2019; Bojnordi Arbab et al., 2024). Prior to the present work, a few studies have examined the angular sizes of AGB stars at sub-mm, mm and longer radio wavelengths. Pertinent results of such studies are summarised in Table 6. To ensure a valid comparison between our measurements of angular sizes and those of earlier studies, we focus on the observations closest to the central frequency of our Band 6 continua (241.75 GHz or 1.24 mm). Vlemmings et al. (2019) studied four nearby oxygen-rich AGB stars (W Hya, Cet (aka Mira A), R Dor, and R Leo) at high angular resolution across multiple wavelength ranges. Higher resolution observations of W Hya (Ohnaka et al., 2024) and R Dor (Vlemmings et al., 2024) are also available. Uniform discs were fit by the authors of these earlier observations, allowing us to directly compare the stellar properties to our sample.
Of these four stars, R Dor is a SRb type variable, while W Hya, Cet and R Leo are all Mira variables. Situating them in our plot of monochromatic luminosity against UD radius in Fig. 6, we find that R Dor sits in the SR quadrant of the plot, with a lower flux and smaller radius than the Mira variables. The three Mira stars lie in the Mira quadrant, upholding the divisions we found. We also considered the pulsation periods of these stars and comparing with our plots of UD properties against period, extending Fig. 7 in Fig. 17. We find that the Mira variables cluster close to the Miras in our sample (with periods of 402 days, 332 days and 310 days for W Hya, Cet, and R Leo, respectively). R Dor has two periods of comparable length to T Mic (332 and 175 days for R Dor, Bedding et al., 1998, and 352 and 178 days for T Mic, Table 1), so sits near T Mic and S Pav on the plot of UD radius against period if we consider the longer period as the primary.
@@
Overall, these past studies agree very well with the trends between radius and period that we see in our sample. Aside from CW Leo, which is most likely an outlier because its size was determined from a Gaussian fit, the Mira variables lie near our trend line (Equation 6), within the uncertainties. Furthermore, when considering the longer period of R Dor, described as Mira-like by Bedding et al. (1998), it also agrees with the trend line (but we note that its smaller radius makes it more typical of the semiregular variables). When using the shorter period for R Dor, the star falls among the semiregular variables in the period-radius plot. There is also good agreement between the literature data and the general trends seen between period and monochromatic luminosity for the ATOMIUM sample. In this case, the main outlier is R Lep, having a larger monochromatic luminosity than other stars with similar periods, but this is probably because the observations were from Band 8 with a higher frequency, making them less directly comparable. We also note that now CW Leo lies close to IRC+10011 and IRC10519, in agreement with our Equation 7 line, which supports our hypothesis that the stars with the longest periods ( days) agree with this trend.
In recent work, Vlemmings et al. (2024) measured the stellar disc of the closest AGB star, R Dor, with ALMA and compared this diameter with the NIR diameter measured by Ohnaka et al. (2019). For the ALMA radius at 225 GHz (1.33 mm, close to our 1.24 mm) the measurements of Vlemmings et al. (2024) and Ohnaka et al. (2019) agree well with our Eq. 4. They report for R Dor , which is the same gradient as in our Eq. 4, but passing through the origin. When we fit to our data but force the line to pass through the origin, we find a slope of 1.38. Both this result and the relation in Eq. 4 are close to being within the uncertainties of the Vlemmings et al. (2024) result for R Dor.